Brief Exercise 9-1

Presented below is information related to Novak Inc.’s inventory.

(per unit)

Skis

Boots

Parkas

Historical cost

$275.50

$153.70

$76.85

Selling price

307.40

210.25

106.94

Cost to sell

27.55

11.60

3.63

Cost to complete

46.40

42.05

30.81


Determine the following: the net realizable value for each item, and the carrying value of each item under LCNRV.

Item Cost NRV LCNRV
Skis $

$

$

Boots

Parkas
Answer #1

Calculation of NRV :-

Skis = 307.40 - 27.55 - 46.40 = 233.45

Boots = 210.25 - 11.60 - 42.05 = 156.60

Parkas = 106.94 - 3.63 - 30.81 = 72.50

Item Cost NRV LCNRV
Skis 275.50 233.45 $233.45
Boots 153.70 156.60 $153.70
Parkas 76.85 72.50 $72.50